OpenStudy (briensmarandache):

sin^4(5x)+cos^4(5x)+2sin^2(5x)cos^2(5x)=1 prove or disprove. can someone help me with this?

OpenStudy (freckles):

Hint a^2+2ab+b^2=(a+b)^2

OpenStudy (solomonzelman):

yes, and we know that: \(\large\color{blue}{ \sin^4(5x)~~~=~~~ \left\{ \sin^2(5x) \right\}^2}\) and \(\large\color{blue}{ \cos^4(5x)~~~=~~~ \left\{ \cos^2(5x) \right\}^2}\)

OpenStudy (briensmarandache):

yeah i did that and got stuck, let me post a pic of my notepad

OpenStudy (solomonzelman):

can you show us what you got after doing that?

OpenStudy (solomonzelman):

I mean, post any work till the point where you got stock. you can draw, type, use latex, or post screenshots.... whatever you prefer.

OpenStudy (solomonzelman):

you forgot about the initial 2sin^2(5x)cos^2(5x) that you had in your problem.

OpenStudy (freckles):

Oh no sin^4(x)+cos^4(x) doesn't equal one

OpenStudy (freckles):

Do you know how to factor a^4+2a^2b^2+b^4?

OpenStudy (briensmarandache):

so then it equals (1-cos^2)^2 assuming where talking about sin^2

OpenStudy (solomonzelman):

What your mistake is that: you did \(\large\color{blue}{ \sin^4(5x)+\cos^4(5x)}\) correctly, but... you had, \(\large\color{blue}{ \sin^4(5x)+\cos^4(5x)\color{red}{+2\sin^2(5x)\cos^2(5x)}=1}\) not just, \(\large\color{blue}{ \cancel{ \sin^4(5x)+\cos^4(5x)=1}}\)

OpenStudy (briensmarandache):

so then it equals (1-cos^2)^2 assuming where talking about sin^2

OpenStudy (briensmarandache):

so what does the first two or three steps of the break down look like? cause im not sure what your saying?

OpenStudy (solomonzelman):

Okay, looking at your equation you can see a \(\large\color{black}{ a^2+2ab+b^2 }\) right? the confusing here is though, that it is disordered. \(\large\color{black}{ a^2+b^2+2ab}\)

OpenStudy (solomonzelman):

See?

OpenStudy (briensmarandache):

and i don't see how this would be factored

OpenStudy (solomonzelman):

\(\large\color{black}{ (a+b)^2=a^2+2ab+b^2}\)

OpenStudy (solomonzelman):

sin^2(5x) is a cos^2(5x) is b

OpenStudy (solomonzelman):

one more spoon, or you're good? :)

OpenStudy (briensmarandache):

ok i got i think

OpenStudy (solomonzelman):

okay, type your work.... or whatever you prefer..

OpenStudy (briensmarandache):

i see the disorder now, cant believe i missed that

OpenStudy (solomonzelman):

But the maun thing that you see it.... :) So go for it:)

OpenStudy (solomonzelman):

oops,, main, not maun. lol

OpenStudy (briensmarandache):

thanks, is it ok if i ping you for any other questions i have? i working on a practice test